Run Planning Meeting Notes, September 3 - 9, 2020

From GlueXWiki
Jump to: navigation, search

<- Back to Run Coordination Meetings:Summer2020

Integrated trigger count

GlueX Statistics Run Sheet

MEDCON 5 restrictions (Bob May email)

  1. adhere to counting room capacity restrictions
  2. maintain 6 ft. separation (briefly passing another person inside 6 ft. while moving about the counting room is inevitable and is OK)
  3. wear masks anytime more than one person present in the counting room
  4. at MEDCON 5 the lab has no cafeteria services because eating "cafeteria style" (eating in common from prepared food) is inconsistent with COVID-19 controls. Staging food in the counting room to share is inconsistent with MEDCON 5 COVID-19 controls.
  5. food and drink consumption - because it requires removing face coverings to consume - should be limited to packaged food and drink brought for personal use that is consumed while maintaining at least 6 ft. from the nearest person and ideally, consumed in a separate room from others.

Wednesday, September 9 at 8:45

  1. Last 24 hours:
    • Continue production runs (3.3B triggers, R73071 - R73082)
    • Beam study (16:50 - 21:20). Works for GEM [1], crystal [2], BCAL voltages [3].
  2. Reports/issues
  3. Run plan
    • Continue production runs

Tuesday, September 8 at 8:30

  1. Last 4 days:
  2. Reports/issues
  3. Run plan
    • Continue production runs
    • Beam study (4h) in the SWING shift (16:30 - 20:30, may vary). Rapid accesses are allowed. Lubomir and Tim are planning to install a small stand at the PS arm. It will take 2-3 h.

Friday, September 4

  1. Last 24 hours:
    • Data taking for DIRC/GEM/TRD finished at 6:40 am. (~4h downtime in the yesterday's evening)
    • De-installation work in progress
    • Also, Sasha is working for crystal. And Nick is working for TAGM in tagger hall.
  2. Reports/issues
    1. Beam position (sharpness of coherent edge)
  3. Run plan
    • Resume production runs after the work in Hall D (and tagger hall). The deinstallation work will be completed at ~10 am (11 am?).
    • Harp scan before starting production runs.

Thursday, September 3

  1. Last 24 hours:
    • -10 am: straight-track data for GEM/TRD
    • Solenoid ramping up during RF recovery, 1352.6 A at 6:40 pm [21].
    • Beam comes back at 9:45 pm. Harp scan.
  2. Reports/issues:
  3. Run plan
    • Continue DIRC commissioning until 6 am tomorrow morning.
    • De-installation work from 6 am (will take 2 - 3 hours)
    • Resume production runs
    • Controlled access to the tagger hall to replace the preamp, should coincide with de-installation work? (Richard, Nick)